# Multicomponent one-dimensional quantum droplets across the mean-field stability regime

### v2

@@ 1 Introduction
The emergence of the self-bound liquid-type configurations known as quantum droplets in contact interacting three-dimensional (3D) bosonic mixtures [24], as well as in dipolar gases [25, 26], is a manifestation of the impact of quantum fluctuations. These may be captured by the Lee-Huang-Yang (LHY) energy [20, 27], representing the first-order correction to the Bogoliubov MF theory, which arrests the expected collapse of the ensuing attractive system at the edge of the MF stability regime [5]. Several properties of these intriguing many-body bound states of matter have already been explored. These refer, exemplarily, to their spectrum [28, 29], collective excitations [20, 28, 30], collisions [31, 30] and coexistence with nonlinear soliton [32, 33] and vortex [34, 35, 36, 37] states, within the context of the LHY theory [22, 38, 21], see also Refs. [39, 40, 41, 42] for the exposition of beyond-LHY effects by deploying ab-initio methods. Experimentally, quantum droplets have been observed in 3D in both homonuclear [24, 43, 44] and heteronuclear [45, 46, 47] mixtures. One of the main challenges faced by these experimental efforts stems from the relatively short droplet lifetimes caused (primarily) by three-body recombination [48]. These lifetimes are increased in relevant 1D settings [49, 50] due to lower densities but also in heteronuclear mixtures [45].
@@ 2.1 Multicomponent bosonic droplet setting
We consider a multicomponent bosonic mixture consisting of $N_{\sigma}$ atoms of mass $m_{\sigma}$ (where $\sigma=A,B,C,D,\dots$) residing in 1D free space and featuring periodic boundary conditions. Such a longitudinal confinement can be experimentally reached with digital micromirror devices [68, 69]. The mixture is assumed to be close to zero temperature, where $s$-wave scattering dominates [70]. Hence, interparticle interactions are represented by contact potentials. Their effective coupling strengths are intracomponent repulsive (i.e. $g_{\sigma\sigma}\equiv g_{\sigma}>0$) and intercomponent attractive ($g_{\sigma\sigma^{\prime}}<0$) or repulsive ($g_{\sigma\sigma^{\prime}}>0$). In a corresponding experiment, they can be adjusted either via Fano-Feshbach resonances [71, 72] through the 3D scattering length or through confinement induced resonances [70] by modifying the transverse trap frequency, $\omega_{\perp}$. This is significantly tight such that transverse excitations are prohibited and the atomic motion is restricted across the elongated $x$-direction as in corresponding 1D experiments [73].
@@ 2.2 Bogoliubov Transformation
This leads to $\mathcal{M}_{b}D=diag(\lambda_{1},...,\lambda_{n},-\lambda_{n+1},...,-\lambda_{2n})$ and therefore the operator $\hat{H}$ takes the form
@@ 4.1 Fully symmetric mixture
The LHY energy of this reduced system, again remains $\sim n^{3/2}$ throughout the MF stability regime $G\in[-g/2,g]$. Moreover, the density independent proportionality factor ranges from $F(G/g)=-\sqrt{6}/\pi$ for $G=-g/2$, to $F(G/g)=-2\sqrt{3}/\pi$ for $G=g$, while it takes its maximum $F(G/g)=-2/\pi$ value at $G=0$, as shown in Fig. 4(a). Interestingly, the strength of quantum fluctuations appears to be larger close to the immiscibility threshold $G=g$ as compared to the one close to the droplet regime $G=-g/2$, as evidenced by the more strongly attractive LHY energy (see e.g. Fig. 4(a) and Eq. (23)). This is in contrast to the two-component mixture where quantum fluctuations contribute equally at the two limits, see in particular Fig. 1(e) and Eq. (13). We attribute this behavior of the three-component system to the aforementioned “asymmetric” MF stability regime.

### v3

@@ 1 Introduction
The emergence of the self-bound liquid-type configurations known as quantum droplets in contact interacting three-dimensional (3D) bosonic mixtures [24, 25], as well as in dipolar gases [26, 27], is a manifestation of the impact of quantum fluctuations. These may be captured by the Lee-Huang-Yang (LHY) energy [20, 28], representing the first-order correction to the Bogoliubov MF theory, which arrests the expected collapse of the ensuing attractive system at the edge of the MF stability regime [5]. Several properties of these intriguing many-body bound states of matter have already been explored. These refer, exemplarily, to their spectrum [29, 30], collective excitations [20, 29, 31], collisions [32, 31] and coexistence with nonlinear soliton [33, 34] and vortex [35, 36, 37, 38] states, within the context of the LHY theory [22, 39, 21], see also Refs. [40, 41, 42, 43] for the exposition of beyond-LHY effects by deploying ab-initio methods. Experimentally, quantum droplets have been observed in 3D in both homonuclear [25, 44, 24] and heteronuclear [45, 46, 47] mixtures. One of the main challenges faced by these experimental efforts stems from the relatively short droplet lifetimes caused (primarily) by three-body recombination [48]. These lifetimes are increased in relevant 1D settings [49, 50] due to lower densities but also in heteronuclear mixtures [45].
@@ 2.1 Multicomponent bosonic droplet setting
We consider a multicomponent bosonic mixture consisting of $N_{\sigma}$ atoms of mass $m_{\sigma}$ (where $\sigma=A,B,C,D,\dots$) residing in 1D free space and featuring periodic boundary conditions. Notice, however, that our results (to be presented below) persist also in the case of a box potential provided that its length is sufficiently large and does not impact the edges of the discussed self-bound droplet configurations. The mixture is assumed to be close to zero temperature, where $s$-wave scattering dominates [68]. Hence, interparticle interactions are represented by contact potentials. Their effective coupling strengths are intracomponent repulsive (i.e. $g_{\sigma\sigma}\equiv g_{\sigma}>0$) and intercomponent attractive ($g_{\sigma\sigma^{\prime}}<0$) or repulsive ($g_{\sigma\sigma^{\prime}}>0$). In a corresponding experiment, they can be adjusted either via Fano-Feshbach resonances [69, 70] through the 3D scattering length or through confinement induced resonances [68] by modifying the transverse trap frequency, $\omega_{\perp}$. This is significantly tight such that transverse excitations are prohibited and the atomic motion is restricted across the elongated $x$-direction as in corresponding 1D experiments [71].
@@ 2.2 Bogoliubov Transformation
This leads to $\mathcal{M}_{b}D={\rm diag}(\lambda_{1},...,\lambda_{n},-\lambda_{n+1},...,-\lambda_{2n})$ and therefore the operator $\hat{H}$ takes the form
@@ 4.1 Fully symmetric mixture
The LHY energy of this reduced system, again remains $\sim n^{3/2}$ throughout the MF stability regime $G\in[-g/2,g]$. Moreover, the density independent proportionality factor ranges from $F(G/g)=-\sqrt{6}/\pi$ for $G=-g/2$, to $F(G/g)=-2\sqrt{3}/\pi$ for $G=g$, while it takes its maximum $F(G/g)=-2/\pi$ value at $G=0$, as shown in Fig. 4(a). Interestingly, the strength of quantum fluctuations appears to be larger close to the immiscibility threshold $G=g$ as compared to the one close to the droplet regime $G=-g/2$, as evidenced by the more strongly attractive LHY energy (see e.g. Fig. 4(a) and Eq. (23)). This is in contrast to the two-component mixture where quantum fluctuations contribute equally at the two limits, see in particular Fig. 1(e) and Eq. (13). We attribute this behavior of the three-component system to the aforementioned “asymmetric” MF stability regime (see also Appendix A for a more elaborated discussion on its asymmetric nature).
@@ Appendix A Details on the stability conditions of the three-component mixture
For the case of the fully symmetric mixture i.e. $g_{A}=g_{B}=g_{C}=g$ and $g_{AB}=g_{AC}=g_{BC}$, the system is significantly simplified and the MF stability requires $-g/2<G<g$. Evidently, the MF stability window becomes reduced compared to the respective two-component one, given by $-g<g_{AB}<g$ as discussed in Sec. 3.1, and asymmetric with respect to the sign of the inter-species interaction strength. An intuitive explanation for this phenomenology is the following. Consider a mixture consisting of $\mathcal{N}$ identical components with $N$ particles per component, featuring intracomponent interactions $g_{\sigma}=g$ and intercomponent ones $g_{\sigma\sigma^{\prime}}=G$, for all $\sigma$, $\sigma^{\prime}$. Ignoring any complications stemming from the exchange statistics and distinguishability, a particle of species $\sigma$ may scatter from $N-1$ $\sigma$-species particles with strength $g$ and $(\mathcal{N}-1)N$ atoms of other components with strength $G$. Hence we expect that, on average, the intracomponent interaction strength will be suppressed by a factor of $\mathcal{N}-1$ (assuming $N\gg 1$) compared to the intercomponent one, yielding an average interaction strength $\propto(G+\frac{g}{\mathcal{N}-1})$. We anticipate the stability of this effective single-component mixture to be ensured as long as its average interaction strength is repulsive (i.e. $G\geq-\frac{g}{\mathcal{N}-1}$), since the effective single-component picture does not impose any limitations at the repulsive side. Indeed this pattern appears to hold for $\mathcal{N}=2,$ (see Sec. 3.1) $\mathcal{N}=3,$ (see Sec. 4.1) and $\mathcal{N}=4$ (see Sec. 5.1), where the lower stability bound corresponds to $G\geq-g$, $G\geq-g/2$, and $G\geq-g/3$ respectively. On the other hand, the only upper bound imposed to the intercomponent interaction strength $G$, stems from the stability of its sub-systems which requires $G<g$ regardless of the number of components. This means that at the repulsive side all components will either be stable and fully overlap or all components will be mutually unstable and phase separate from all other components. A hypothetical partially phase separated configuration, with $\mathcal{M}$ overlapping components in one cluster and $\mathcal{N}-\mathcal{M}$ in the other, would feature average intra-cluster interaction strengths $\propto G+\frac{g}{\mathcal{M}-1}$ and $\propto G+\frac{g}{\mathcal{N}-\mathcal{M}-1}$ respectively, both of which are larger than the inter-cluster coupling strength $G$ (considering only repulsive interactions). As such, there is no reason for the clusters to phase separate, which explains why such configurations do not seem to appear.